# Volumetric Surfaces: Representing Fuzzy Geometries with Multiple Meshes

### v1

@@ Abstract
High-quality real-time view synthesis methods are based on volume rendering, splatting, or surface rendering. While surface-based methods generally are the fastest, they cannot faithfully model fuzzy geometry like hair. In turn, alpha-blending techniques excel at representing fuzzy materials but require an unbounded number of samples per ray (P1). Further overheads are induced by empty space skipping in volume rendering (P2) and sorting input primitives in splatting (P3). These problems are exacerbated on low-performance graphics hardware, e.g. on mobile devices. We present a novel representation for real-time view synthesis where the (P1) number of sampling locations is small and bounded, (P2) sampling locations are efficiently found via rasterization, and (P3) rendering is sorting-free. We achieve this by representing objects as semi-transparent multi-layer meshes, rendered in fixed layer order from outermost to innermost. We model mesh layers as SDF shells with optimal spacing learned during training. After baking, we fit UV textures to the corresponding meshes. We show that our method can represent challenging fuzzy objects while achieving higher frame rates than volume-based and splatting-based methods on low-end and mobile devices. Project page: s-esposito.github.io/volsurfs.
Figure 1: Our Volumetric Surfaces representation (b) consists of $k$ lightweight, semi-transparent mesh shells that efficiently render fuzzy geometries with a limited number of samples $(3\!\leq\!k\!\leq\!9)$ via rasterization. Our image quality surpasses that of surface-based methods (a) and approaches the quality of 3D Gaussian splatting (c). Unlike splatting methods, our sorting-free representation enables faster rendering, particularly on low-end or mobile graphics hardware.
@@ 1 Introduction
Real-time rendering is hard on mobile devices due to limited processing power, memory, and thermal constraints. Recent methods for real-time view synthesis can be categorized according to their rendering paradigm. On the one hand, we have surface-based methods like BakedSDF [51] or BOG [34], where rendering a pixel requires only reading appearance data from a single sampling location along the ray. On the other hand, we have volume-based methods like SMERF [8] or 3DGS [17], where rendering a pixel requires reading data from multiple sample locations along the ray. As a result, surface-based methods are generally faster than volume-based techniques, which struggle to achieve interactive frame rates on mobile devices [34]. While recent surface-based methods are also capable of representing thin structures like individual strands of grass [34, 55], they still lag behind volume-based methods, especially when it comes to reconstructing highly fuzzy geometry like hair or plush. Even if possible from a reconstruction perspective, a purely surface-based representation might be too memory-inefficient for representing extremely fuzzy objects [2]. Towards our goal of real-time view synthesis of fuzzy objects on mobile devices, we therefore focus on finding a more efficient volumetric formulation.
A crucial factor for the efficiency of a volumetric representation is the number of samples needed along a ray. For state-of-the-art methods like SMERF or 3DGS, tens or hundreds of samples per ray are often required. Also the intrinsics of the respective rendering algorithm influence performance. In volume rendering, we need to traverse an additional data structure to skip empty space, which requires additional memory bandwidth and leads to suboptimal thread coherence. For splatting, we need to sort primitives according to their distance from the camera, which is hard to implement efficiently on platforms with limited GPGPU capabilities. To address these problems of existing approaches, we design a novel representation, where the number of sampling points per ray is bounded to a small number (three to nine). Our final representation consists of multiple semi-transparent layers wrapped around the reconstructed object. Our representation is trained such that we can rasterize layers in a fixed order from outermost to innermost. Hence, we neither need to resort to expensive empty space skipping nor sorting during rendering.
@@
However, it is non-trivial to learn the optimal spacing between individual layers; we tackle this by representing them as separate SDFs during training. Before training all layers, we start by training a single opaque SDF that prevents degenerate solutions. We then add the additional layers, constraining them from intersecting one another. To further increase the expressivity of our representation, while keeping a low number of layers, we make each layer’s transparency depend on the viewing direction. All layers are optimized toward smooth solutions, so that each can be baked into a lightweight mesh for efficient hardware-accelerated rasterization. The simplicity of our meshes enables computing a high-quality UV parameterization, which is often problematic for highly complex, monolithic meshes [34]. Finally, we optimize UV textures of spherical harmonics (SH) coefficients on the fixed geometry defined by our meshes.
@@
We demonstrate how our approach renders significantly faster than volume-based methods, enabling high frame rates on a wide range of commodity devices, while simultaneously being more capable at representing fuzzy objects than single-surface approaches.
@@ 2.1 Real-time View Synthesis
Neural radiance fields (NeRFs) achieved an impressive leap of quality by fitting a 3D scene representation via differentiable volume rendering to multi-view images [24]. NeRFs represent the scene implicitly as a multi-layer perceptron (MLP) [23, 30, 4]. Evaluating an MLP is arithmetically intensive, leading to slow inference. To overcome this, a number of works explore faster representations based on 3D grids [20, 32, 14, 53, 11, 9], triplanes [3, 33, 8], hash grids [26, 40], or explicit primitives [1, 48, 36, 6, 17]. 3D Gaussians have gained traction lately due to their fast training and rendering [17]. While these representations greatly increase efficiency, rendering a pixel still requires a potentially high number of samples per ray.
(a) Volumetric
(b) Surface
(c) Ours.
Figure 2: Different sampling strategies: a) volumetric rendering’s dense sampling; b) single sampling point, as in surface rendering; c) our method, only sampling the first intersection with each surface.
@@
DoNeRF reduces sample count by only sampling around depth values predicted by a separate MLP, which, however, requires access to ground-truth depth maps [27]. AdaNeRF explicitly introduces sampling sparsity during the course of training [18]. HybridNeRF introduces a hybrid surface–volume representation that encourages surfaces over volumetric rendering [42]. AdaptiveShells restrict sampling to a small shell around the surface [47]. This shell is represented by a triangle mesh, which is rasterized to find the sampling range. Similar to them, we also use triangle rasterization to find sampling locations. However, our method finds all sampling locations via rasterization, while AdaptiveShells still employs standard volume rendering within the surface envelope. Consequently, AdaptiveShells still represents the scene with a 3D volume texture, while our multi-layer formulation enables the usage of 2D textures. Our method also differs in how we obtain the surface envelope. While AdaptiveShells learns a single SDF and determines the width of the envelope with a spatially-varying kernel, we progressively learn multiple, separate SDFs.
NRDF employs a two-layered mesh [44]. Both the outer and inner mesh are rasterized and feature vectors are read from the surfaces, which are decoded by a convolutional neural network (CNN). In turn, our textures directly store view-dependent opacities and colors and we do not require a CNN decoder. Further, NRDF heuristically extracts the two layers via marching cubes with different density thresholds, while we propose a progressive multi-SDF training strategy for obtaining layered meshes. Similar to us, QuadFields also aggregate opacities and view-dependent colors from multiple ray-triangle intersections [38]. However, they require blending in depth order, necessitating ray casting or depth peeling, whereas our formulation allows us to simply blend layers in a fixed order. Simultaneously, our meshes are significantly more compact than theirs.
@@
Another line of work aims for a single sample per ray. MobileNeRF represents the scene with a coarse proxy mesh equipped with a binary opacity mask to increase expressivity [5]. BakedSDF [51], NeRF2Mesh [41], NeRFMeshing [31] and BOG[34] first perform a 3D reconstruction of the scene, convert their respective training-time representation into a mesh, and then fit an appearance model to the mesh. While these single-surface approaches are fastest, they struggle with reconstructing fuzzy geometries. Even if it would be possible to accurately reconstruct fuzzy geometry from multi-view data, representing individual hairs with hard geometry requires a lot of memory [2].
2.2 3D Reconstruction
@@
Earlier methods for 3D reconstruction were often based on image matching [10, 37]. More recent works directly fit level-set representations via differentiable rendering [28, 49, 29, 43]. To improve convergence, many approaches convert an SDF to a volumetric density on-the-fly, which is then used for standard volume rendering [50, 45, 46, 54, 35, 19]. A number of recent methods also show that depth maps rendered from 2D or 3D Gaussians can be converted into a high-quality surface representation via volumetric fusion [15, 7, 55, 56].
@@ 3 Preliminaries
In this section we provide a brief introduction to volume- and surface-based representations and related notations.
Figure 3: Our $k$-SDF network, predicting $k$ distance values as described in Section 4.1.
Rendering equation: A ray $\mathbf{r}$ in 3D space is parameterized by its origin $\mathbf{o}\in\mathbb{R}^{3}$ and unit direction $\mathbf{v}\in\mathbb{S}^{2}$. A 3D point at distance $t$ along ray $r$ is given by $\mathbf{r}(t)=\mathbf{o}+\mathbf{v}t$. Let $\sigma$ denote a density field, which assigns a nonnegative density value $\sigma(\mathbf{x})\in\mathbb{R}_{+}$ to each 3D point $\mathbf{x}\in\mathbb{R}^{3}$, and let $\xi$ be a vector field providing an RGB color $\xi(\mathbf{x},\mathbf{v})$ for each point $\mathbf{x}\in\mathbb{R}^{3}$ and direction $\mathbf{v}\in\mathbb{S}^{2}$. We can render $\xi$ along ray $r$ for a given density field $\sigma$ using the following equation [24]:
@@
Volume-based representations (NeRF): NeRF models a scene as a volume – with absorption and emission but without scattering effects [39] – parametrized as $(\sigma,\xi)$. NeRF numerically approximates Equation 1 with quadrature [22], densely sampling rays to render each pixel, as illustrated in Figure 2(a).
@@
Here, the spread of densities around the surface (zero-level set) is controlled by the scalar $\beta$. A small $\beta$ results in fuzzy densities, while in the limit $\beta\rightarrow\infty$ densities are sharp, impulse functions for points on the implicit surface. While NeuS regards $\beta$ as a learnable parameter, [35] showed how controlling it explicitly leads to better reconstructions. In our case, scheduling $\beta$ ensures that densities shift from being widely spread to being peaked by the end of the training. When densities are peaked, the representation can be baked into a mesh suitable for efficient rendering. However, in this case, all appearance information is condensed on a single point (Figure 2(b)); therefore, surface-based methods are not able to model semi-transparent surfaces, strongly limiting their ability to handle mixed pixels often representing thin structures, which are notoriously missed [51].
@@ 4 Method
In this section, we describe our proposed representation, Volumetric Surfaces, in detail. We begin with its implicit geometry definition, then cover rendering, training, and baking. We also discuss and justify our design choices.
@@ 4.1 $k$-SDF
Our new representation, called $k$-SDF, models $k$ surfaces as distinct signed distance fields $\{d_{1},\ldots,d_{k}\}$. To composite the surfaces, $k$-SDF is endowed with an additional transparency field $\alpha$, which assigns a view-dependent transparency value $\alpha(\mathbf{x},\mathbf{v})\in[0,1]$ to each 3D point $\mathbf{x}$. The transparency field enables modeling semi-transparent surfaces. Additionally, we have a color field $\xi$ to render images. Given a ray $\mathbf{r}$, we can render a $k$-SDF $(d_{1:k},\alpha,\xi)$ using the following rendering equation that generalizes Equation 1:
$\mathcal{R}_{\beta}(\mathbf{r}\mid d_{1:k},\alpha,\xi)=\sum_{i=1}^{k}\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\xi)\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\alpha)w_{i}\text{,}$ (3)
@@
where $w_{i}=\prod_{j=1}^{i}\mathcal{(}1-\mathcal{R}_{\beta}(\mathbf{r}\mid d_{j},\alpha))$ and, for the sake of compactness, $\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\gamma)$ stands for $\mathcal{R}(\mathbf{r}\mid\phi_{\beta}\circ d_{i},\gamma)$. Similarly to other SDF-based representations [45], our per-surface density field is derived from the signed distance field $d_{i}$ via a parametrized logistic function (Equation 2).
@@
Support surfaces as shells: Blending weights in Equation 3 are order-dependent, but we cannot guarantee that lower-ranked densities are positioned closer to the camera. To avoid sorting, we model our set of surfaces as shells. This ensures that layers are traversed in ray-intersection order. We model our $k$-SDF with a main surface represented as an SDF $d$, and a set of $k-1$ support surfaces represented as offset fields $\{o_{2},\ldots,o_{k}\}$ (Figure 4) from $d$’s zero-level set. The signed distance functions for each surface are thus given by:
$d_{1:k}$ $=(d,d+o_{2},\ldots,d+o_{k})\,\text{.}$ (4)
A surface defined by a positive offset is contained inside the main one, while a negative offset yields a surface containing the main one. We force the sign of a predicted offset using a softplus activation function multiplied by the desired sign. To model more than two support surfaces, while enforcing the order property, we run a cumulative sum (separately on negative and positive offsets) and use resulting displacements to parameterize the support surfaces. Figure 3 illustrates a simplified version of our model, where all offsets are positive.
Rendering individual surfaces: We sample $n$ points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ for each ray $\mathbf{r}$. RGB and transparency fields ($\xi$, $\alpha$) are conditioned on the encoded sample positions $\psi(\mathbf{x}_{i})$, view direction $\mathbf{v}$, the per-sample, per-surface SDF normal $\mathbf{n}_{i,j}$ and a geometric feature vector $\mathbf{z}_{i}$ predicted by the shared MLP decoder of the $k$-SDF net. The rendered color $\mathcal{C}_{j}(\mathbf{r})$ and transparency $\mathcal{A}_{j}(\mathbf{r})$ for each surface $j$ and ray $\mathbf{r}$ are then given by:
$\mathcal{C}_{j}(\mathbf{r})$ $=\sum_{i=1}^{n}w_{i,j}\xi(\psi(\mathbf{x}_{i}),\mathbf{v},\mathbf{n}_{i,j},\mathbf{z}_{i})\text{,}$ (5)
$\mathcal{A}_{j}(\mathbf{r})$ $=\sum_{i=1}^{n}w_{i,j}\alpha(\psi(\mathbf{x}_{i}),\mathbf{v},\mathbf{n}_{i,j},\mathbf{z}_{i})\text{.}$ (6)
[…]
4.2 Surface Blending

### v2

# Volumetric Surfaces: Representing Fuzzy Geometries with Layered Meshes

@@ Abstract
High-quality view synthesis relies on volume rendering, splatting, or surface rendering. While surface rendering is typically the fastest, it struggles to accurately model fuzzy geometry like hair. In turn, alpha-blending techniques excel at representing fuzzy materials but require an unbounded number of samples per ray (P1). Further overheads are induced by empty space skipping in volume rendering (P2) and sorting input primitives in splatting (P3). We present a novel representation for real-time view synthesis where the (P1) number of sampling locations is small and bounded, (P2) sampling locations are efficiently found via rasterization, and (P3) rendering is sorting-free. We achieve this by representing objects as semi-transparent multi-layer meshes rendered in a fixed order. First, we model surface layers as signed distance function (SDF) shells with optimal spacing learned during training. Then, we bake them as meshes and fit UV textures. Unlike single-surface methods, our multi-layer representation effectively models fuzzy objects. In contrast to volume and splatting-based methods, our approach enables real-time rendering on low-power laptops and smartphones.
a) PermutoSDF
b) Ours (7-Mesh)
c) 3DGS
d) Ours (7-Mesh)
Figure 1: Our Volumetric Surfaces representation (b) consists of $k$ lightweight, semi-transparent mesh shells that efficiently render fuzzy geometries with a limited number of samples $(3\!\leq\!k\!\leq\!9)$ via rasterization. Our image quality surpasses that of surface-based methods (a) and approaches the quality of 3D Gaussian splatting (c). Unlike splatting methods, our sorting-free representation enables faster rendering, particularly on low-power or mobile graphics hardware. Project page: https://autonomousvision.github.io/volsurfs.
@@ 1 Introduction
Real-time rendering on mobile devices is challenging due to limited processing power, memory, and thermal constraints. Recent methods for real-time view synthesis can be categorized according to their rendering paradigm. On the one hand, we have surface-based methods like BakedSDF [60] or BOG [43], where rendering a pixel requires only reading appearance data from a single sampling location along the ray. On the other hand, we have volume-based methods like SMERF [10] or 3DGS [24], where rendering a pixel requires reading data from multiple sample locations along the ray. As a result, surface-based methods are generally faster than volume-based techniques, which struggle to achieve interactive frame rates on mobile devices [43]. While recent surface-based methods are also capable of representing thin structures like individual strands of grass [43, 64], they still lag behind volume-based methods, especially when it comes to reconstructing highly fuzzy geometry like hair or plush. Even if possible from a reconstruction perspective, a purely surface-based representation might be too memory-inefficient for representing extremely fuzzy objects [4]. Towards our goal of real-time view synthesis of fuzzy objects on mobile devices, we therefore focus on finding a more efficient volumetric formulation. A key factor in the efficiency of volumetric representations is the number of samples required along a ray. State-of-the-art methods, such as SMERF and 3DGS, often require tens or even hundreds of samples per ray. Additionally, the intrinsic characteristics of the rendering algorithm impact performance. In volume rendering, traversing an extra data structure to skip empty space increases memory bandwidth usage and results in suboptimal thread coherence. For splatting, primitives must be sorted by their distance from the camera, a task that is challenging to implement efficiently on platforms with limited GPGPU capabilities. For these reasons, such approaches are not suited for current mobile hardware, e.g. low-cost smartphones. To address these limitations, we propose a differentiable representation that bounds the number of sampling points per ray to a small range (three to nine) and is sorting-free. This allows us, unlike 3DGS or SMERF, to achieve real-time rendering on low-cost smartphones.
@@
Textured shells [27, 26, 23] have long been used in computer graphics to simulate fuzzy surfaces while minimizing geometric complexity. These are modeled as concentric, uniformly spaced, semi-transparent layers, and are still widely used in modern game production. Inspired by these techniques, our representation learns adaptively spaced layers around a reconstructed object via gradient-based optimization. This allows for rasterization in a fixed order, from outermost to innermost, without the need for expensive empty space skipping or sorting during rendering. However, it is non-trivial to learn the optimal spacing between individual layers; we tackle this by representing them as separate signed distance functions (SDFs) during training. Before training all layers, we start by training a single opaque SDF that prevents degenerate solutions. We then add the additional layers, constraining them from intersecting one another. To further increase the expressivity of our representation, while keeping a low number of layers, we make each layer’s transparency depend on the viewing direction. All layers are optimized toward smooth solutions so that each can be baked into a lightweight mesh for efficient hardware-accelerated rasterization. The simplicity of our meshes enables computing a high-quality UV parameterization, which is often problematic for highly complex, monolithic meshes [43]. Finally, we optimize UV textures of spherical harmonics (SH) coefficients on the fixed geometry defined by our meshes.
@@
We demonstrate how our approach renders significantly faster than volume-based and splatting-based methods, enabling high frame rates on a wide range of commodity devices, while simultaneously being more capable at representing fuzzy objects than single-surface approaches.
@@ 2 Related Work
Real-Time View Synthesis: Neural radiance fields (NeRFs) achieved an impressive leap of quality by fitting a 3D scene representation via differentiable volume rendering to multi-view images [34]. NeRFs represent the scene implicitly as a multi-layer perceptron (MLP) [33, 39, 6]. Evaluating an MLP is arithmetically intensive, leading to slow inference. To overcome this, a number of works explore faster representations based on 3D grids [29, 41, 19, 62, 14, 12], triplanes [5, 42, 10], hash grids [35, 49], or explicit primitives [2, 57, 45, 8, 24]. 3D Gaussians have gained traction lately due to their fast training and rendering [24]. While these representations enhanced efficiency, rendering each pixel may still require a large number of samples per ray.
@@
DONeRF reduces sample count by only sampling around depth values predicted by a separate MLP, which, however, requires access to ground-truth depth maps [36]. AdaNeRF explicitly introduces sampling sparsity during the course of training [25]. HybridNeRF introduces a hybrid surface–volume representation that encourages surfaces over volumetric rendering [51]. AdaptiveShells restrict sampling to a small shell around the surface [56]. This shell is represented by a triangle mesh, rasterized to find the sampling range. Unlike AdaptiveShells, which uses volume rendering, our method finds all sampling locations via rasterization, enabling the use of 2D textures instead of 3D volume textures. Additionally, while AdaptiveShells learns a single SDF with a spatially-varying kernel, we learn multiple SDFs. NDRF [53] extracts two mesh layers via marching cubes with varying density thresholds and rasterizes them to feature vectors that are decoded by a convolutional neural network (CNN). In turn, we store view-dependent opacities and colors as SH textures. Similar to us, QuadFields also aggregate opacities and view-dependent colors from multiple ray-triangle intersections [47]. However, these methods require significantly larger meshes, intersection calculations for all triangles along a ray, and a costly sorting process. In contrast, our formulation enables blending in a fixed order.
@@
Another line of work aims for a single sample per ray. MobileNeRF [7] represents the scene with a coarse proxy mesh equipped with a binary opacity mask to increase expressivity. BakedSDF [60], NeRF2Mesh [50], NeRFMeshing [40] and BOG [43] first perform a 3D reconstruction of the scene, convert their respective training-time representation into a mesh, and then fit an appearance model to the mesh. While these single-surface approaches are faster, they struggle to reconstruct fuzzy geometries. Even with accurate reconstruction from multi-view data, representing individual hairs explicitly would require substantial memory [4].
Gaussian frosting [18] anchors Gaussians on a tight surface shell to enhance editability. Like our approach, GaussianShellMaps [1] utilize a multi-layer mesh representation. However, unlike our approach, GaussianShellMaps uses fixed shells rather than adaptively learning them, and predicts per-shell Gaussians. Further these methods still fall short of achieving real-time rendering on budget smartphones due to their reliance on splatting techniques.
@@
3D Reconstruction: Earlier methods for 3D reconstruction were often based on image matching [13, 46]. More recent works directly fit level-set representations via differentiable rendering [37, 58, 38, 52]. To improve convergence, many approaches convert an SDF to a volumetric density on-the-fly, which is then used for standard volume rendering [59, 54, 55, 63, 44, 28]. Recent methods demonstrate converting 2D or 3D Gaussians into high-quality surface representations, either through volumetric fusion [20, 9, 64, 65], or using a Poisson reconstruction algorithm [17].
@@ 3 Preliminaries
In this section, we provide a brief introduction to volume- and surface-based representations and related notations.
Rendering Equation: A ray $\mathbf{r}$ in 3D space is parameterized by its origin $\mathbf{o}\in\mathbb{R}^{3}$ and unit direction $\mathbf{v}\in\mathbb{S}^{2}$. A 3D point at distance $t$ along ray $\mathbf{r}$ is given by $\mathbf{r}(t)=\mathbf{o}+\mathbf{v}t$. Let $\sigma$ denote a density field, which assigns a nonnegative density value $\sigma(\mathbf{x})\in\mathbb{R}_{+}$ to each 3D point $\mathbf{x}\in\mathbb{R}^{3}$, and let $\boldsymbol{\xi}$ be a vector field providing an RGB color $\xi(\mathbf{x},\mathbf{v})$ for each point $\mathbf{x}\in\mathbb{R}^{3}$ and direction $\mathbf{v}\in\mathbb{S}^{2}$. We can render $\boldsymbol{\xi}$ along ray $\mathbf{r}$ for a given density field $\sigma$ using the following equation [34]:
@@
Volume-Based Representations (NeRF): NeRF models a scene as a volume – with absorption and emission but without scattering effects [48] – parametrized as $(\sigma,\boldsymbol{\xi})$. NeRF numerically approximates Equation 1 with quadrature [32], densely sampling rays to render each pixel (Figure 1a).
a Volumetric
b Surface
c Ours
Figure 2: Sampling strategies: ($a$) volumetric rendering’s dense sampling; ($b$) single sampling point, as in surface rendering; ($c$) our method, only sampling the first intersection with each surface.
@@
Here, the spread of densities around the surface (zero-level set) is controlled by the scalar $\beta$. A small $\beta$ results in fuzzy densities, while in the limit $\beta\rightarrow\infty$ densities are sharp impulse functions for points on the implicit surface. While NeuS regards $\beta$ as a learnable parameter, Rosu and Behnke [44] showed how controlling it explicitly leads to better reconstructions. In our case, scheduling $\beta$ ensures that densities shift from being widely spread to being peaked by the end of the training. When densities are peaked, the representation can be baked into a mesh suitable for efficient rendering. However, in this case, all appearance information is condensed on a single point (Figure 1b); therefore, surface-based methods are not able to model semi-transparent surfaces, strongly limiting their ability to handle mixed pixels often representing thin structures, which are notoriously missed [60].
@@ 4 Method
In this section, we describe our proposed representation, Volumetric Surfaces, in detail. We begin with its implicit geometry definition, then cover rendering, training, and baking. We also discuss and justify our design choices. Comprehensive architectural details can be found in the supplement.
a
b
Figure 3: ($a$) High-level architecture of our $k$-SDF network, predicting $k$ distance values as described in Section 4.1. For simplicity of visualization, all offsets are positive. We highlight trainable components. For additional architectural details, refer to the supplementary material. ($b$) 1D example visualization of the output $d_{1:k}$ when evaluating the network at a sample point $\mathbf{x}$ along a ray. Signed distances are shown as solid lines, while $\beta$-controlled integration weights are represented as dotted lines.
@@
Our new representation, called $k$-SDF, models $k$ surfaces as distinct signed distance fields $\{d_{1},\ldots,d_{k}\}$. To composite the surfaces, $k$-SDF is endowed with an additional transparency field $\alpha$, which assigns a view-dependent transparency value $\alpha(\mathbf{x},\mathbf{v})\in[0,1]$ to each 3D point $\mathbf{x}$. The transparency field enables modeling semi-transparent surfaces. We generalize Equation 1 to render a $k$-SDF $(d_{1:k},\alpha,\boldsymbol{\xi})$ as:
$\mathcal{R}_{\beta}(\mathbf{r}\mid d_{1:k},\alpha,\boldsymbol{\xi})=\sum_{i=1}^{k}\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\boldsymbol{\xi})\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\alpha)w_{i}\text{,}$ (3)
@@
where $w_{i}=\prod_{j=1}^{i}\mathcal{(}1-\mathcal{R}_{\beta}(\mathbf{r}\mid d_{j},\alpha))$ and, for the sake of compactness, $\mathcal{R}_{\beta}(\mathbf{r}\mid d_{i},\gamma)$ stands for $\mathcal{R}(\mathbf{r}\mid\phi_{\beta}\circ d_{i},\gamma)$. Our per-surface density field is derived from the signed distance field $d_{i}$ as in Equation 2.
@@
Support Surfaces as Shells: Blending weights in Equation 3 are order-dependent, but lower-ranked densities are not necessarily positioned closer to the camera. To avoid sorting, we model the set of surfaces as shells, ensuring that layers are traversed in ray-intersection order. We model our $k$-SDF with a main surface represented as an SDF $d$, and a set of $k\!-\!1$ support surfaces represented as offset fields $\{o_{2},\ldots,o_{k}\}$ (Figure 4) from $d$’s zero-level set. The signed distance functions for each surface are thus given by: $d_{1:k}=(d,d+o_{2},\ldots,d+o_{k})$. A surface defined by a positive offset is contained inside the main one, while a negative offset yields a surface containing the main one. To model multiple support surfaces while enforcing order, we perform a cumulative sum over predicted relative offsets $\{\hat{o}_{2},\ldots,\hat{o}_{k}\}$ (separately for negative and positive offsets) and use the resulting absolute displacements $\{o_{2},\ldots,o_{k}\}$ to parameterize the surfaces. Figure 3 illustrates our model.
Rendering Individual Surfaces: Equation 1 is approximated with quadrature [34] and evaluated at $n$ discrete points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ sampled along a ray $\mathbf{r}$. For surface $j$, we render its color $\mathcal{C}_{j}(\mathbf{r})$ and transparency $\mathcal{A}_{j}(\mathbf{r})$ as:
$\mathcal{C}_{j}(\mathbf{r})$ $=\sum_{i=1}^{n}w_{i,j}\boldsymbol{\xi}(\mathbf{x}_{i},\mathbf{v},\mathbf{n}_{i,j},\mathbf{z}_{i})\text{,}$ (4)
$\mathcal{A}_{j}(\mathbf{r})$ $=\sum_{i=1}^{n}w_{i,j}\alpha(\mathbf{x}_{i},\mathbf{v},\mathbf{n}_{i,j},\mathbf{z}_{i})\text{,}$ (5)
where RGB and transparency fields ($\boldsymbol{\xi}$, $\alpha$) are conditioned on sample position $\mathbf{x}_{i}$, view direction $\mathbf{v}$, SDF normal $\mathbf{n}_{i,j}$, and feature vector $\mathbf{z}_{i}$, an additional output of the $k$-SDF model. Sample weights $w_{i,j}$ are computed as in Wang et al. [54].
4.2 Surfaces Blending